OpenStudy (anonymous):

Please help! simplify. x^2+4x-21 over x+7 times x-3 over x^2-5x+6

OpenStudy (anonymous):

=(x^2+(7x-3x)-21)/(x+7) =(x^2+7x)-(3x+21)/(x+7) =x(x+7)-3(x+7)/(x+7) =(x-3)(x+7)/(x+7) =(x-3) (1) now we do the 2nd part (x-3)/(x^2-5x+6) =(x-3)/(x^2-3x-2x+6) =(x-3)/(x(x-3)-2(x-3)) =(x-3)/((x-3)(x-2) =1/(x-2) (2) now multiplying 1 nd 2 (x-3) /(x-2)...... ans

OpenStudy (anonymous):

is it helped

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

u welkum :P